OpenStudy (anonymous):

Also, solve for x: log(7x+1)-log(x^2+13)=0 I honestly cannot solve this.

OpenStudy (anonymous):

x=4 or 3

OpenStudy (anonymous):

log(x^2+13)= log (7x+1) x^2+13 = 7x+1 x² - 7x + 12 = 0 => x = 3 , x = 4

OpenStudy (campbell_st):

take the exponential to get rid of the logs then 7x + 1 = x^2 + 13 (x -3)(x-4) = 0 x = 3,4

OpenStudy (anonymous):

Oooh see I thought it was log (7x+1)/(x^2+13) Thanks!!